# Time Resolved Absorption of Six Chemical Species With MAROON-X Points to a Strong Drag in the Ultra Hot Jupiter TOI-1518 b

### v3

@@ 4.2 Injection and Cross-Correlation maps
Inject the spectra at each orbital phase into the three components that were removed during the data reduction process, using the velocity value from Equation 1.
@@
Figs. 11 and 12 show the CCF signals of and that we obtain for each of the models, with the real data plotted on top.
@@ 4.3 Discussion
Fig.11 (Left Panel) shows the Fe trails of the five GCMs simulations, along with the Doppler shifts of measured in section 3.2. All models show a signal that is blueshifting with time. This behaviour was previously observed by Ehrenreich et al. (2020); Borsa et al. (2021) and discussed in many papers (Wardenier et al., 2021, 2023; Savel et al., 2022; Beltz et al., 2022, 2023). Because the planet rotates during the transit, the start of the transit probes the dayside of the leading limb and the nightside of the trailing limb. It is thus dominated by the hotter leading limb. The planet rotation red-shifts the leading limb by 3.5 km/s, which is compensated by the blue shift of the day-to-night winds. In the second part of the transit, the opposite happens and the signal is dominated by the trailing limb, which is blueshifted by rotation. The change from being dominated by the leading and the trailing limb during transit is responsible for the blueshifting of the signal with time.
@@
Furthermore, the global shift of the planetary trace is linked to the day-to-night winds that blueshift both leading and trailing hemisphere signals. These day-to-night winds are strongly affected by drag. As shown in Fig.11, models without drag produce an iron signal that is too blueshifted compared to the observations. Overall, the planetary signal stands between the stronger drag models ( and s). When looking at the amplitude and at the FWHM of the signal, however, we see that the s model is favoured, because the s provides a signal that is too small and not wide enough compared to data.
If we now look at the trace (Fig.12), we see that the s overestimates the blue shift, whereas the s ( in purple) is a much better match. For the amplitude of the model is unaffected by the drag and the FHWM errorbars are too large to differentiate between the two strong drag models. The apparent contradiction between the Fe trace favouring the s model and the trace favouring the stronger, s drag model could be an indication of the presence of ohmic drag in the atmosphere. Indeed, as shown by Beltz et al. (2022), Ohmic drag strength strongly depends on planetary location, being stronger in the hotter dayside and at lower pressures. Because the equilibrium is temperature dependent, the signal naturally probes more into the hotter planetary dayside than the Fe signal, where the temperatures are hotter and magnetic drag effects are expected to be larger.
Fig.13 shows the trails for the 6 species of Fig.7 with the signals of the stronger drag models. For , , , and the signal straddles between both these GCMs, as in the case of Fe. We therefore confirm that the strong drag scenario is preferred by the signal of all the other species. A similar agreement with an intense drag model atmosphere was already observed in the case of WASP-121 b by Wardenier et al. 2024, pointing out that strong drag is common in ultra-hot Jupiter atmospheres.
The trail of the is significantly different from the ones of the other species, as it starts strongly red-shifted. This is clearly not captured by the model. We calculated the contribution functions of our model, and find that the CCF of most species (including Ca, Fe and ) probe the and bar layers whereas the probes layers up to . We fitted directly the resolved individual lines using a method similar to that of Prinoth et al. (2024) and found a transit depth corresponding to an effective radius of . This value is similar to the planetary Roche radius (R at Lagrange Point 1 following Gu et al. (2003)). As a consequence, we believe that the lines likely probe the outflow of the planet. This region shows a different atmospheric flow than the deeper regions, with an important contribution from planet-star interactions, not modeled in our GCM. Our results for are similar to the recent observations of H-alpha for WASP-121b (Seidel et al., 2025).
@@ 5.2 Results of the retrieval analysis
The mass prior is set using the SOPHIE observation results (see Table 7) as a free parameter that follows a Gaussian distribution with Mp = 1.83 M. For this analysis, we consider the two detectors independently as for the CCF, and we add the likelihood from the red and the blue detector analysis and then from both transit. For the retrieval analysis, we used three number of principal components as for the CCF analysis. As shown before, the PCA does not strongly affect the planetary lines, apart from the ones of the . However, because the retrieval considers all species at the same time, we could not have different PCA numbers per species. The results of the measured abundances are summarized in Fig. 14 and the whole corner plot is given in the Appendix.(Fig. 26).
@@ 5.3 Discussion of the retrieval analysis
The Kp and Vres distribution of the retrieval analysis(Fig. 26) are consistent with the iron signal detected in section 3.2. The measured Vres () and Kp () also agree with the detection of Cabot et al. (2021) ( km/s and Kp= km/s). We observed that the velocity parameters obtained are influenced predominantly by the strongest absorber, particularly Fe.
@@
Our model retrieves an iron abundance of . This corresponds to 0.07 to 1.62 solar enrichment in iron. This remains consistent with a solar (and then stellar) metallicity. We have checked that the iron abundance was robust to different assumptions in the retrieval, and that it was not changing when different other species were added.
@@
The retrieval results for the other chemical species are more surprising and highlight the difficulty of atmospheric retrievals for transit spectroscopy at high-spectral resolution. First, the ionized species, such as and have retrieved abundances that are much higher than solar. These high abundances, up to are unlikely in a gas giant atmosphere. Instead, we believe that these could stem from a lack of flexibility in the forward model. Indeed, ionized species are likely to probe different parts of the atmosphere (both vertically and longitudinally) than neutral species, because the spatial repartition of neutral and ionized forms of the same species is anti-correlated. This can lead to signals at different temperatures and Kp and Vres (e.g. Table 3). However, given that the temperature, Kp and Vres for all species is determined by the neutral iron lines, which represent the strongest signal, the abundances of the other species are likely biased to compensate for the Kp/Vres offset.
@@
For lines with a strong Kp/Vsys shift and small opacities, such as , , or , the retrieval leads only to upper limits, which are all consistent with a solar enrichment.
@@
For the neutral calcium and magnesium lines, the retrieval provides a detection with an abundance of and , corresponding to 0.0002 to 0.034 times the solar abundance for and 0.0004 to 0.339 times solar for . The retrieved value for neutral calcium is significantly lower than that of iron, and it is, for now, unclear whether this is due to a physical effect (e.g. most being into ) or a bias in the retrieval. Finally, our results for deserve a special mention.
@@
Despite using the latest line list, the multi-species retrieval did not successfully detect . To explore this issue further, we conducted a single-species retrieval focused solely on (see Fig. 27). This approach resulted in a detection with a temperature around 2700 K, compared to 3500 K in the multi-species retrieval, along with a different distribution (Fig. 15). The measured value of in such a framework is , corresponding to an enrichment of 1.10 to 45708.81 times solar. The discrepancy in the distribution resembles the patterns observed in the map (see Table 4). This difference could be attributed to either dynamical factors—if probes specific regions of the atmosphere—or inaccuracies in the positioning of the lines within the line list. This situation underscores the limitations of the current framework in accurately deriving abundance for species that are settled in different distribution.
@@
Further observations with bluer or redder instruments to detect molecules or ionized species are needed to determine why such elements are subsolar. Moreover, using a non-isothermal profile instead would have a dual effect on the spectrum. First, it would produce a similar effect to an abundance gradient by stretching the spectral lines. Second, it would enhance the abundance gradient in the case of chemical equilibrium, further impacting the observed spectral features. Incorporating more realistic temperature and abundance gradients in future retrieval analyses will be essential to better constrain the abundances of ionized species.
@@ 6 Conclusions
Our cross-correlation analysis focused on detecting atomic and molecular species within the atmosphere. We report the detection of 14 different species. High-resolution spectroscopy allowed us to identify ionized and neutral species with ionized metals such as , , and . The detection of these ionized species, as opposed to their neutral counterparts, underscores the extreme temperatures of TOI-1518 b’s atmosphere, where thermal ionization is significant. Additionally, we reported the detection of vanadium oxide (), a critical absorber that has implications for thermal inversions in ultra-hot Jupiters. The presence of VO adds a crucial piece to the puzzle of understanding the thermal structure and chemical processes in such extreme environments. This was possible thanks to the newer HyVO line list that should be used in other studies were detection of VO failed with previous line lists.
We investigated the atmospheric wind dynamics by analyzing the blueshift of multiple species in the observed spectra as a function of the planet’s orbital phase. The blueshift of iron is consistent with previous observations of other ultra-hot Jupiters. By combining the signal from different species, particularly and Fe, and comparing them with global circulation models , we conclude that a strong drag is needed ( to explain the data. seem to need more substantial drag compared to Fe, which can be due to the increased Ohmic drag in the planetary dayside. Finally, we find that the trail of is very different from the other species, which is likely due to probing the escaping atmosphere above the Roche lobe while all other species probe deeper atmospheric pressures.
The retrieval analysis provided constraints on the abundance of various chemical species in the atmosphere. We report an abundance of iron of , which is consistent with a solar enrichment. The retrieval being driven mainly by iron, we believe that the iron abundances are robust. However, most other species are detected at a different /, which our retrieval framework does not consider. As a consequence, we believe that the retrieved abundance for the other species are strongly biased. We study in more detail this bias for the specific case of , for which our main retrieval does not find a signal, but where a retrieval without iron lead to a detection at a different temperature, , and an abundance of 1.10 to 45708.81 times solar.
@@ Appendix A Measurement of the planetary mass
The FWHM of the Gaussian is unknown, but is expected to be small by comparison to the FWHM of the rotational profile. We attempted free and fixed values for that Gaussian FWHM, which did not significantly modify our results. We finally fixed the FWHM of the Gaussian profile to 7 km/s on each exposure, which is a common value for such kind of stars. The FWHM of the rotational profil was free to vary among exposures, and we obtained values corresponding to km/s, with a dispersion of km/s. This is in good agreement with the values km/s reported by Cabot et al. 2021. The measured RVs (see Table 6) are derived from the fitted center of the Gaussian and rotational profiles, constrained to be identical for a given epoch.

### v4

@@ 4.2 Injection and cross-correlation maps
Inject the spectra at each orbital phase into the three components that were removed during the data reduction process, using the velocity value from Eq. 1.
@@
Figs. 11 and 12 show the CCF signals of Fe and Fe+ that we obtain for each of the models, with the real data plotted on top.
@@ 4.3 Discussion
Fig.11 (Left Panel) shows the Fe trails of the five GCMs simulations, along with the Doppler shifts of Fe measured in Sect. 3.2. All models show a signal that is blueshifting with time. This behavior was previously observed by Ehrenreich et al. (2020); Borsa et al. (2021) and discussed in many papers (Wardenier et al., 2021, 2023; Savel et al., 2022; Beltz et al., 2022, 2023). Because the planet rotates during the transit, the start of the transit probes the dayside of the leading limb and the nightside of the trailing limb. It is thus dominated by the hotter leading limb. The planet rotation red-shifts the leading limb by 3.5 km/s, which is compensated by the blue shift of the day-to-night winds. In the second part of the transit, the opposite happens and the signal is dominated by the trailing limb, which is blueshifted by rotation. The change from being dominated by the leading and the trailing limb during transit is responsible for the blueshifting of the signal with time.
@@
Furthermore, the global shift of the planetary trace is linked to the day-to-night winds that blueshift both leading and trailing hemisphere signals. These day-to-night winds are strongly affected by drag. As shown in Fig.11, models without drag produce an iron signal that is too blueshifted compared to the observations. Overall, the planetary signal stands between the stronger drag models ( and s). When looking at the amplitude and at the FWHM of the signal, however, we see that the s model is favored, because the s provides a signal that is too small and not wide enough compared to data.
If we now look at the Fe+ trace (Fig.12), we see that the s overestimates the blue shift, whereas the s ( in purple) is a much better match. For Fe+ the amplitude of the model is unaffected by the drag and the FHWM errorbars are too large to differentiate between the two strong drag models. The apparent contradiction between the Fe trace favoring the s model and the Fe+ trace favoring the stronger, s drag model could be an indication of the presence of ohmic drag in the atmosphere. Indeed, as shown by Beltz et al. (2022), Ohmic drag strength strongly depends on planetary location, being stronger in the hotter dayside and at lower pressures. Because the Fe/Fe+ equilibrium is temperature dependent, the Fe+ signal naturally probes more into the hotter planetary dayside than the Fe signal, where the temperatures are hotter and magnetic drag effects are expected to be larger.
Fig.13 shows the trails for the 6 species of Fig.7 with the signals of the stronger drag models. For Ca, Na, Mg, and Ca+ the signal straddles between both these GCMs, as in the case of Fe. We therefore confirm that the strong drag scenario is preferred by the signal of all the other species. A similar agreement with an intense drag model atmosphere was already observed in the case of WASP-121 b by Wardenier et al. 2024, pointing out that strong drag is common in ultra-hot Jupiter atmospheres.
The trail of the Ca+ is significantly different from the ones of the other species, as it starts strongly red-shifted. This is clearly not captured by the model. We calculated the contribution functions of our model, and find that the CCF of most species (including Ca, Fe and Fe+) probe the and bar layers whereas the Ca+ probes layers up to . We fitted directly the resolved individual Ca+ lines using a method similar to that of Prinoth et al. (2024) and found a transit depth corresponding to an effective radius of . This value is similar to the planetary Roche radius (R at Lagrange Point 1 following Gu et al. (2003)). As a consequence, we believe that the Ca+ lines likely probe the outflow of the planet. This region shows a different atmospheric flow than the deeper regions, with an important contribution from planet-star interactions, not modeled in our GCM. Our results for Ca+ are similar to the recent observations of H-alpha for WASP-121b (Seidel et al., 2025).
@@ 5.2 Results of the retrieval analysis
The mass prior is set using the SOPHIE observation results (see Table 7) as a free parameter that follows a Gaussian distribution with Mp = 1.83 M. For this analysis, we consider the two detectors independently as for the CCF, and we add the likelihood from the red and the blue detector analysis and then from both transit. For the retrieval analysis, we used three number of principal components as for the CCF analysis. As shown before, the PCA does not strongly affect the planetary lines, apart from the ones of the Ca+. However, because the retrieval considers all species at the same time, we could not have different PCA numbers per species. The results of the measured abundances are summarized in Fig. 14 and the whole corner plot is given in the Appendix.(Fig. 26).
@@ 5.3 Discussion of the retrieval analysis
The Kp and Vres distribution of the retrieval analysis(Fig. 26) are consistent with the iron signal detected in Sect. 3.2. The measured Vres ( km/s) and Kp ( km/s) also agree with the detection of Cabot et al. (2021) ( km/s and Kp= km/s). With a of approximately -25 km/s compared to the derived from the SOPHIE data, and a Vres around -3 km/s, the retrieval is also capturing the blueshift of the signal as a function of the planet’s orbital phase, as detailed in Sect. 4. We observed that the velocity parameters obtained are influenced predominantly by the strongest absorber, particularly Fe.
@@
Our model retrieves an iron abundance of Fe=. This corresponds to 0.07 to 1.62 solar enrichment in iron. This remains consistent with a solar (and then stellar) metallicity. We have checked that the iron abundance was robust to different assumptions in the retrieval, and that it was not changing when different other species were added.
@@
The retrieval results for the other chemical species are more surprising and highlight the difficulty of atmospheric retrievals for transit spectroscopy at high-spectral resolution. First, the ionized species, such as Fe+ and Ca+ have retrieved abundances that are much higher than solar. These high abundances, up to Fe= are unlikely in a gas giant atmosphere. Instead, we believe that these could stem from a lack of flexibility in the forward model. Indeed, ionized species are likely to probe different parts of the atmosphere (both vertically and longitudinally) than neutral species, because the spatial repartition of neutral and ionized forms of the same species is anti-correlated. This can lead to signals at different temperatures and Kp and Vres (e.g. Table 3). However, given that the temperature, Kp and Vres for all species is determined by the neutral iron lines, which represent the strongest signal, the abundances of the other species are likely biased to compensate for the Kp/Vres offset.
@@
For lines with a strong Kp/Vsys shift and small opacities, such as Cr, V, Ti or VO, the retrieval leads only to upper limits, which are all consistent with a solar enrichment.
@@
For the neutral calcium and magnesium lines, the retrieval provides a detection with an abundance of Ca= and Mg=, corresponding to 0.0002 to 0.034 times the solar abundance for Ca and 0.0004 to 0.339 times solar for Mg. The retrieved value for neutral calcium is significantly lower than that of iron, and it is, for now, unclear whether this is due to a physical effect (e.g. most Ca being into Ca+) or a bias in the retrieval. Finally, our results for VO deserve a special mention.
@@
Despite using the latest VO line list, the multi-species retrieval did not successfully detect VO. To explore this issue further, we conducted a single-species retrieval focused solely on VO (see Fig. 27). This approach resulted in a detection with a temperature around 2700 K, compared to 3500 K in the multi-species retrieval, along with a different distribution (Fig. 15). The measured value of VO in such a framework is VO, corresponding to an enrichment of 1.10 to 45708.81 times solar. The discrepancy in the distribution resembles the patterns observed in the map (see Table 4). This difference could be attributed to either dynamical factors—if VO probes specific regions of the atmosphere—or inaccuracies in the positioning of the lines within the VO line list. This situation underscores the limitations of the current framework in accurately deriving abundance for species that are settled in different distribution.
@@
Further observations with bluer or redder instruments to detect molecules or ionized species are needed to determine why such elements are subsolar. Moreover, using a non-isothermal profile instead would have a dual effect on the spectrum. First, it would produce a similar effect to an abundance gradient by stretching the spectral lines. Second, it would enhance the abundance gradient in the case of chemical equilibrium, further impacting the observed spectral features. Incorporating more realistic temperature and abundance gradients in future retrieval analyzes will be essential to better constrain the abundances of ionized species.
@@ 6 Conclusions
Our cross-correlation analysis focused on detecting atomic and molecular species within the atmosphere. We report the detection of 14 different species. High-resolution spectroscopy allowed us to identify ionized and neutral species with ionized metals such as Fe+, Ca+, and Ti+. The detection of these ionized species, as opposed to their neutral counterparts, underscores the extreme temperatures of TOI-1518 b’s atmosphere, where thermal ionization is significant. Additionally, we reported the detection of vanadium oxide (VO), a critical absorber that has implications for thermal inversions in ultra-hot Jupiters. The presence of VO adds a crucial piece to the puzzle of understanding the thermal structure and chemical processes in such extreme environments. This was possible thanks to the newer HyVO line list that should be used in other studies were detection of VO failed with previous line lists.
We investigated the atmospheric wind dynamics by analyzing the blueshift of multiple species in the observed spectra as a function of the planet’s orbital phase. The blueshift of iron is consistent with previous observations of other ultra-hot Jupiters. By combining the signal from different species, particularly Fe+ and Fe, and comparing them with GCMs , we conclude that a strong drag is needed ( to explain the data. Fe+ seem to need more substantial drag compared to Fe, which can be due to the increased Ohmic drag in the planetary dayside. Finally, we find that the trail of Ca+ is very different from the other species, which is likely due to Ca+ probing the escaping atmosphere above the Roche lobe while all other species probe deeper atmospheric pressures.
The retrieval analysis provided constraints on the abundance of various chemical species in the atmosphere. We report an abundance of iron of Fe, which is consistent with a solar enrichment. The retrieval being driven mainly by iron, we believe that the iron abundances are robust. However, most other species are detected at a different /, which our retrieval framework does not consider. As a consequence, we believe that the retrieved abundance for the other species are strongly biased. We study in more detail this bias for the specific case of VO, for which our main retrieval does not find a signal, but where a retrieval without iron lead to a detection at a different temperature, , and an abundance of 1.10 to 45708.81 times solar.
@@ Appendix A Measurement of the planetary mass
The FWHM of the Gaussian is unknown, but is expected to be small by comparison to the FWHM of the rotational profile. We attempted free and fixed values for that Gaussian FWHM, which did not significantly modify our results. We finally fixed the FWHM of the Gaussian profile to 7 km/s on each exposure, which is a common value for such kind of stars. The FWHM of the rotational profile was free to vary among exposures, and we obtained values corresponding to km/s, with a dispersion of km/s. This is in good agreement with the values km/s reported by Cabot et al. 2021. The measured RVs (see Table 6) are derived from the fitted center of the Gaussian and rotational profiles, constrained to be identical for a given epoch.